the company has a 70 million dollar outstanding bond issue that matures in 5 years with a semi annual coupon of 25.00 and selling price of 841.74. The company also has 5,000,000 shares of common stock selling at 19.00. The company's beta is 1.13 the t-bill rate is 2.6% and the market rate is 7.4%. the company preferred stock pays 1.45 dividend and has a price of 29.00 a share. The total company value is $200,000,000

What is WACC?
Answer #1

MV of debt = par value*selling price/1000

=70*841.74/1000=58.9218m

MV of equity = price*shares = 5*19=95m

MV of preferred equity =total company value -MV of debt-MV of equity = 200-95-58.9218

=46.0782

Weight of debt = Value of debt/Total Assets Value
= 58.9218/200
=0.2946
Weight of equity = Value of equity/Total Assets Value
= 95/200
=0.475
Weight of Preferred equity = Value of Preferred equity/Total Assets Value
= 46.0782/200
=0.2304
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 2.6 + 1.13 * (7.4 - 2.6)
Cost of equity% = 8.02
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =5x2
841.74 =∑ [(5*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^5x2
                   k=1
YTM = 9.0001580038
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 9.0001580038*(1-0)
= 9.0001580038
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 1.45/29*100
=5
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=9*0.2946+8.02*0.475+5*0.2304
WACC% = 7.61